Can you please sovle me this equation 2y+3x=1050
Y+x=450

Answer by mananth(16946)

2y+3x=1050
y+x=450
multiply by -2
-2y-2x=-900
2y+3x=1050
Add this equation to the first one
x= 150
x+y =450
Plug x
y+150=450
y=450-150
y=300